Hello! I am working on my Algebra Homework and I am stuck on these 2 problems.. Can you please help me Tutor?
1) Give 3 equivalent fractions to 3/5. I could only think of 9/15!
2) Write in standard form: fifteen thousand, four and 53ten- thousandth.
Please Reply :) Thank You!
-Lauren

Answer by checkley77(12844)

3/5*3=9/15*2=18/30*5=90/150*3=270/450 etc.